# Soft-Label Caching and Sharpening for Communication-Efficient Federated Distillation

### v1

@@ III-B SCARLET: System Design and Workflow
Algorithm 1 describes the overall workflow of SCARLET. While the workflow follows that of DS-FL [3], SCARLET introduces two major differences. First, after local training, each client selectively transmits soft-labels only for public dataset samples that are not currently cached on the server, thereby avoiding the redundant transmissions that are typical in previous methods. This caching mechanism, detailed in Section III-C, synchronizes soft-label storage between the server and clients, significantly reducing communication overhead compared to DS-FL, which retransmits soft-labels in every round.
Second, SCARLET introduces a novel aggregation process at the server after collecting soft-labels. We propose Enhanced ERA, an adaptive aggregation strategy that enables robust control over the entropy of aggregated predictions. In contrast to the original ERA used in DS-FL, which often requires careful tuning of the temperature parameter, Enhanced ERA adopts a new parameterization that allows flexible and stable adjustment of label sharpness. This design enhances robustness and stability under varying degrees of data heterogeneity across clients. Enhanced ERA is described in detail in Section III-D.
Together, the soft-label caching mechanism and Enhanced ERA distinctly position SCARLET as a highly communication-efficient and robust alternative to existing distillation-based FL methods. The subsequent subsections provide detailed descriptions and analysis of these components.
@@ III-C Soft-Label Caching in SCARLET
Figure 2: Illustration of the soft-label caching mechanism in SCARLET. (Left) When public data is not cached, clients transmit their soft-labels to the server, where they are aggregated using Enhanced ERA and stored in the server cache before synchronizing with client caches. (Right) When public data is cached, both the server and clients retrieve soft-labels from their respective caches for knowledge distillation, eliminating the need for additional soft-label communication.
@@
function UpdateGlobalCache($\mathcal{C}_{g}$, $\hat{z}^{t}$, $\mathcal{I}^{t}$, $t$, $D$)
     Input: Cache $\mathcal{C}_{g}$, soft-labels $\hat{z}^{t}$, indices $\mathcal{I}^{t}$,
@@
     Output: Cache signals $\gamma^{t}$
     Initialize $\gamma^{t}\leftarrow[\;]$
     for all $i,z\in(\mathcal{I}^{t},\hat{z}_{t})$ do
         if $i\notin\text{dom}(\mathcal{C}_{g})$ then
              $\mathcal{C}_{g}\leftarrow\mathcal{C}_{g}[i\mapsto(z,t)]$
              Append NEWLY_CACHED to $\gamma^{t}$
         else
              $(z_{c},t_{c})\leftarrow\mathcal{C}_{g}(i)$
              if $t-t_{c}\leq D$ then
                  Append CACHED to $\gamma^{t}$
              else
                  $\mathcal{C}_{g}\leftarrow\mathcal{C}_{g}\setminus\{i\}$
                  Append EXPIRED to $\gamma^{t}$
              end if
         end if
     end for
     return $\gamma^{t}$
end function
function UpdateLocalCache($\mathcal{C}_{k}$, $\gamma^{t-1}$, $\hat{z}_{\text{req}}^{t-1}$, $\mathcal{I}^{t-1}$)
     Input: Cache $\mathcal{C}_{k}$, cache signals $\gamma^{t-1}$,
@@
     Output: Soft-labels $\hat{z}^{t-1}$
     Initialize $\hat{z}^{t-1}\leftarrow[\;]$
     Initialize queue $Q\leftarrow\hat{z}_{\mathrm{req}}^{t-1}$triangleright treat as FIFO queue
     for all $i,z\in(\mathcal{I}^{t-1},\hat{z}^{t})$ do
         if $\gamma^{t-1}[i]$ is NEWLY_CACHED then
              $\mathcal{C}_{k}\leftarrow\mathcal{C}_{k}[i\mapsto Q.\text{pop}()]$
              Append $\mathcal{C}_{k}(i)$ to $\hat{z}^{t-1}$
         else if $\gamma^{t-1}[i]$ is CACHED then
              Append $\mathcal{C}_{k}(i)$ to $\hat{z}^{t-1}$
         else if $\gamma^{t-1}[i]$ is EXPIRED then
              $\mathcal{C}_{k}\leftarrow\mathcal{C}_{k}\setminus\{i\}$
              Append $Q.\text{pop}()$ to $\hat{z}^{t-1}$
         end if
     end for
     return $\hat{z}^{t-1}$
end function
Mechanism Description:  The soft-label caching mechanism is introduced in SCARLET to address the communication inefficiency caused by redundant transmission of soft-labels in distillation-based FL methods. Specifically, SCARLET reduces communication overhead by caching and reusing soft-labels across communication rounds.
@@
The caching mechanism introduces minimal additional costs compared to the communication overhead it reduces. Each round, the server sends cache synchronization signals, denoted as $\gamma^{t}$ in Algorithm 2, to clients. These signals indicate the cache status of public dataset samples and can be represented as integers, resulting in an additional communication cost of $O(\lvert\mathcal{P}^{t}\rvert)$, where $\lvert\mathcal{P}^{t}\rvert$ is the number of public dataset samples selected per round. For memory, both the server and clients require $O(\lvert\mathcal{P}\rvert\times N)$ to cache soft-labels, where $\lvert\mathcal{P}\rvert$ is the total size of the public dataset and $N$ is the number of classes. This memory usage is small compared to model sizes in FL and can be further optimized by storing caches on disk if necessary.
@@
Preliminary Evaluation of Sample Cache Ratio:  Figure 3 illustrates the relationship between the cache duration $D$ and the proportion of cached samples per round, based on a simulation using the public dataset $\mathcal{P}$. In each communication round, 10% of the samples in $\mathcal{P}$ were randomly selected to form $\mathcal{P}^{t}$, and if a selected sample was already cached, its cached soft-label was reused instead of being updated.
@@
As $D$ increases, a greater number of selected samples remain valid in the cache across rounds, leading to a higher cache hit rate and a corresponding reduction in the need to retransmit soft-labels. However, excessively large values of $D$ reduce the frequency of cache updates, increasing the likelihood of reusing outdated soft-labels—particularly those that significantly deviate from the true labels—which may negatively impact training.
These results demonstrate that the caching module effectively reduces communication overhead while providing a tunable trade-off between the reuse of cached soft-labels and their freshness, controlled via the cache duration parameter $D$.
The soft-label caching mechanism serves as a versatile and scalable enhancement for FL systems, offering a general-purpose solution to reduce communication costs in distillation-based FL methods. Its design is independent of the underlying aggregation or distillation mechanism, making it easily compatible with existing methods. By integrating this module, other approaches can achieve significant communication efficiency improvements with minimal modifications to their core algorithms.
III-D Enhanced ERA in SCARLET
@@
(c) Entropy vs. $T$ using ERA for (a) and (b), shown on the left and right, respectively
@@
(d) Entropy vs. $\beta$ using Enhanced ERA for (a) and (b), shown on the left and right, respectively
Figure 4: Comparison of ERA and Enhanced ERA with different parameter settings. Panels (a) and (b) show the original high-entropy (low-confidence) and low-entropy (high-confidence) soft-labels, respectively. Panel (c) illustrates how ERA rapidly reduces entropy by decreasing temperature $T$, particularly for (a) compared to (b). Panel (d) demonstrates how Enhanced ERA provides smoother and more flexible entropy control by varying $\beta$, ensuring that $\beta=1$ recovers the original entropy for both (a) and (b).
In [3], the ERA method was introduced to adjust aggregated soft-labels, mitigating training stagnation caused by noisy predictions. ERA employs a temperature-adjusted softmax function to systematically reduce the entropy of aggregated client predictions. However, excessively reducing entropy can negatively impact model accuracy. Furthermore, as demonstrated in our preliminary experiments (Fig. 4(c)), ERA exhibits sensitivity to the temperature $T$, with optimal values significantly dependent on the initial entropy of the soft-labels. This sensitivity complicates parameter tuning.
To address this limitation, SCARLET proposes Enhanced ERA, defined as:
@@
where $\overline{z^{t}}=\frac{1}{K}\sum_{k=1}^{K}z_{k}^{t}$ represents the averaged local soft-labels, and $\beta$ controls aggregation sharpness. As $\beta$ increases beyond 1, Enhanced ERA gradually sharpens the aggregated soft-labels. This sharpening effect is formally described using the concept of majorization: the resulting distribution for a larger $\beta$ is majorized by the distribution for a smaller $\beta$. This majorization relationship guarantees a monotonic reduction in entropy due to the Schur-concavity of Shannon entropy. For completeness, we provide a proof in the Appendix A.
Enhanced ERA utilizes the Power Mean function for aggregating client predictions, resulting in smoother and more consistent entropy control. Figure 4(d) illustrates the relationship between entropy and the hyperparameter $\beta$ when applying Enhanced ERA. Importantly, when $\beta=1$, Enhanced ERA preserves the original entropy irrespective of the initial entropy level, making it particularly suitable for low-entropy soft-labels that require no entropy reduction. Additionally, Enhanced ERA provides gradual entropy reduction for high-entropy soft-labels. While the achievable entropy reduction is more moderate compared to the original ERA (unless extreme $\beta$ values are employed), the experiments in the following section demonstrate that even this moderate reduction substantially improves model accuracy in distillation-based FL scenarios.
By providing controlled entropy adjustments, Enhanced ERA enables SCARLET to effectively handle different degrees of non-IID data distributions. Specifically, $\beta=1$ maintains the original entropy for weak non-IID scenarios, while setting $\beta>1$ allows gradual sharpening of soft-labels for stronger non-IID conditions. This controlled adjustment simplifies parameter tuning and leads to robust performance across diverse federated learning scenarios. Optimal tuning of $\beta$ itself, however, remains an open research question and is considered future work. This paper primarily focuses on introducing the fundamental mechanisms of soft-label caching and Enhanced ERA, with detailed discussions regarding hyperparameter tuning provided in Section V.

### v3

@@ III-B SCARLET: System Design and Workflow
Algorithm 1 describes the overall workflow of SCARLET. At the beginning of each communication round $t$, the server selects a subset of the public dataset, $\mathcal{P}^{t}$, to be used for distillation.
While the SCARLET framework can accommodate both full and partial client participation, we first assume the fundamental case of full client participation (i.e., all $K$ clients participate in every round) for analytical clarity. This full participation setting allows a focused analysis on the introduction of our two primary contributions: the soft-label caching mechanism (Section III-C and Algorithm 2) and the Enhanced ERA mechanism (Section III-E).
The necessary considerations for the more practical and complex scenario of partial client participation (i.e., where clients participate intermittently), which extends this fundamental case, will be detailed separately in Section III-D. The subsequent experimental evaluation in Section IV-B will then validate SCARLET’s effectiveness under both of these settings: the full participation scenario (Sections IV-B2 to IV-B5) and the partial participation scenario (Section IV-B6).
While the workflow follows that of the baseline DS-FL [dsfl], SCARLET introduces two major differences.
First, after local training, each client selectively transmits soft-labels only for public dataset samples that are not currently cached on the server. This selective transmission is based on the request list $\mathcal{I}_{\text{req}}^{t}$ computed by the server (Algorithm 1, line 10). This is managed by a global cache $\mathcal{C}_{g}$ on the server, which stores previously aggregated soft-labels. This caching mechanism, detailed in Section III-C, synchronizes soft-label storage between the server and clients via cache signals $\gamma^{t}$ (generated in line 41 and sent in line 14), significantly reducing communication overhead compared to DS-FL, which retransmits soft-labels in every round.
Second, SCARLET introduces a novel aggregation process at the server after collecting soft-labels (Algorithm 1, line 35). We propose Enhanced ERA, an adaptive aggregation strategy that enables robust control over the entropy of aggregated predictions. In contrast to the conventional ERA used in DS-FL, which often requires careful tuning of the temperature parameter, Enhanced ERA adopts a new parameterization that allows flexible and stable adjustment of label sharpness. This design enhances robustness and stability under varying degrees of data heterogeneity across clients. Enhanced ERA is described in detail in Section III-E.
Together, the soft-label caching mechanism and Enhanced ERA distinctly position SCARLET as a highly communication-efficient and robust alternative to existing distillation-based FL methods. The subsequent sections provide detailed descriptions and analysis of these components.
@@ III-C Soft-Label Caching in SCARLET
(a) Public data (not cached)
(b) Public data (cached)
Figure 2: Illustration of the soft-label caching mechanism in SCARLET. When public data is not cached (2a), clients transmit their soft-labels to the server, where they are aggregated using Enhanced ERA and stored in the server cache before synchronizing with client caches. When public data is cached (2b), both the server and clients retrieve soft-labels from their respective caches for knowledge distillation, eliminating the need for additional soft-label communication.
@@
1:function UpdateGlobalCache($\mathcal{C}_{g}$, $\hat{z}^{t}$, $\mathcal{I}^{t}$, $t$, $D$)
2:  Input: Cache $\mathcal{C}_{g}$, soft-labels $\hat{z}^{t}$, indices $\mathcal{I}^{t}$,
@@
3:  Output: Cache signals $\gamma^{t}$
4:  Initialize $\gamma^{t}\leftarrow[\;]$
5:  for all $i,z\in(\mathcal{I}^{t},\hat{z}_{t})$ do
6:   if $\mathcal{C}_{g}(i)\text{ does not exist}$ then
7:     $\mathcal{C}_{g}(i)\leftarrow(z,t)$
8:     Append NEWLY_CACHED to $\gamma^{t}$
9:   else
10:     $(z_{c},t_{c})\leftarrow\mathcal{C}_{g}(i)$
11:     if $t-t_{c}\leq D$ then
12:      Append CACHED to $\gamma^{t}$
13:     else
14:      Delete $\mathcal{C}_{g}(i)$
15:      Append EXPIRED to $\gamma^{t}$
16:     end if
17:   end if
18:  end for
19:  return $\gamma^{t}$
20:end function
21:
22:function UpdateLocalCache($\mathcal{C}_{k}$, $\gamma^{t-1}$, $\hat{z}_{\text{req}}^{t-1}$, $\mathcal{I}^{t-1}$)
23:  Input: Cache $\mathcal{C}_{k}$, cache signals $\gamma^{t-1}$,
@@
24:  Output: Soft-labels $\hat{z}^{t-1}$
25:  Initialize $\hat{z}^{t-1}\leftarrow[\;]$
26:  Initialize queue $Q\leftarrow\hat{z}_{\mathrm{req}}^{t-1}$ $\triangleright$ treat as FIFO queue
27:  for all $i,z\in(\mathcal{I}^{t-1},\hat{z}^{t})$ do
28:   if $\gamma^{t-1}[i]$ is NEWLY_CACHED then
29:     $\mathcal{C}_{k}(i)\leftarrow Q.\text{pop}()$
30:     Append $\mathcal{C}_{k}(i)$ to $\hat{z}^{t-1}$
31:   else if $\gamma^{t-1}[i]$ is CACHED then
32:     Append $\mathcal{C}_{k}(i)$ to $\hat{z}^{t-1}$
33:   else if $\gamma^{t-1}[i]$ is EXPIRED then
34:     Delete $\mathcal{C}_{k}(i)$
35:     Append $Q.\text{pop}()$ to $\hat{z}^{t-1}$
36:   end if
37:  end for
38:  return $\hat{z}^{t-1}$
39:end function
The soft-label caching mechanism is introduced in SCARLET to address the communication inefficiency caused by redundant transmission of soft-labels in distillation-based FL methods. Specifically, SCARLET reduces communication overhead by caching and reusing soft-labels across communication rounds.
@@
[…]
@@
A key advantage of this mechanism is that the impact of $D$ can be estimated using a lightweight, standalone simulation without executing the full FL process. This simulation, detailed in Appendix A, models only the random sampling process—not the FL training itself—to rapidly predict the cache hit rate over time for a given $D$. Fig. 3 shows the result of such a simulation, modeling a scenario where 10% of the total public dataset $\mathcal{P}$ is randomly selected as $\mathcal{P}^{t}$ each round. The figure confirms that $D$ dynamically controls the ratio of cached samples. Due to the random sampling, this ratio fluctuates before decaying and stabilizing over the rounds.
This simulation reveals that $D$ directly governs the cache’s refresh frequency. In distillation-based FL, the transmission of soft-labels for these public samples constitutes the majority of communication cost. Therefore, even a conservative duration, such as $D<100$, results in a high proportion of public data samples being cached, offering significant communication savings. However, for $D\geq 200$, the simulation shows periods where the cache ratio approaches 1.0. During these rounds, all samples used for distillation are retrieved from the cache. This means the system is training on identical, outdated soft-labels, and communication is reduced to only the cache signals. This risks both accuracy degradation from stale knowledge and wasted communication rounds.
@@
This analysis of the simulation results clarifies the fundamental trade-off: As $D$ increases, a greater number of selected samples remain valid in the cache across rounds, leading to a higher cache hit rate and a corresponding reduction in the need to retransmit soft-labels. However, as the simulation highlights, excessively large values of $D$ reduce the frequency of cache updates, increasing the likelihood of reusing outdated soft-labels—particularly those that significantly deviate from the true labels—-which may negatively impact training. These results demonstrate that the caching module effectively reduces communication overhead while providing a tunable trade-off between the reuse of cached soft-labels and their freshness, controlled via the cache duration parameter $D$.
This lightweight simulation thus provides a valuable tool for identifying a reasonable $D$ that balances cost reduction with cache freshness. The practical guideline for selecting $D$ using this simulation procedure, as well as its effectiveness, is discussed in the final paragraph of Section IV-B4.
III-D Handling Partial Client Participation
When clients participate intermittently, the integrity of the local cache ($\mathcal{C}_{k}$) becomes a critical challenge. A client rejoining a session after being offline for several rounds will hold a stale cache, which is not synchronized with the server’s global cache ($\mathcal{C}_{g}$). This client cannot correctly reconstruct the full soft-label set $\hat{z}^{t-1}$ from the standard update package ($\gamma^{t-1},\hat{z}_{req}^{t-1}$) as described in Algorithm 2.
To address this, we introduce a “catch-up package” mechanism. When a client $k$ is selected for round $t$, the server checks if the client participated in the previous round t1t-1. If the client is synchronized, having participated in t1t-1, it receives the standard, lightweight update package as described in Algorithm 1 (line 14), containing only the cache signals $\gamma^{t-1}$ and the soft-labels for newly computed samples $\hat{z}_{req}^{t-1}$.
Conversely, if the client is stale, meaning it did not participate in round t1t-1, its local cache is not synchronized with the latest updates. Therefore, in addition to the standard lightweight package for round t1t-1, the server sends a catch-up package. This supplementary package contains the aggregated soft-labels required for the client to update its cache for all the rounds it was inactive. Before performing the distillation step, the client first uses this catch-up package to update its local cache $\mathcal{C}_{k}$ for all indices in $\mathcal{I}^{t-1}$, ensuring it is fully synchronized. It then proceeds with the distillation step (Algorithm 1, line 22) using the now-complete cache information.
This mechanism guarantees that all participating clients possess the correct teacher soft-labels for distillation. On the downlink, this incurs an additional communication cost specifically for stale clients, corresponding to the differential updates accumulated since their last participation to synchronize with the server’s latest cache, thereby avoiding the need to synchronize all clients in every round. In contrast, the uplink contribution remains consistently communication-efficient: clients strictly compute and transmit soft-labels $z_{k}^{t}$ only for the samples in the server’s request list $\mathcal{I}_{req}^{t}$ (Algorithm 1, line 31).
III-E Enhanced ERA in SCARLET
@@
(c) Entropy vs. $T$ using ERA for (4a) and (4b); shown left and right, respectively
@@
(d) Entropy vs. $\beta$ using Enhanced ERA for (4a) and (4b); shown left and right, respectively
Figure 4: Comparison of ERA and Enhanced ERA with different parameter settings. (4a) and (4b) show the original high-entropy (low-confidence) and low-entropy (high-confidence) soft-labels, respectively. (4c) illustrates how ERA rapidly reduces entropy by decreasing temperature $T$, particularly for (4a) compared to (4b). (4d) demonstrates how Enhanced ERA provides smoother and more flexible entropy control by varying $\beta$, ensuring that $\beta=1$ recovers the original entropy for both (4a) and (4b).
A primary challenge in distillation-based FL arises when aggregating soft-labels from clients with non-IID data. Simply averaging their heterogeneous predictions often results in an ambiguous global soft-label with high entropy, which provides a poor training signal and degrades performance.
To address this, [dsfl] introduced Entropy Reduction Aggregation (ERA), defined in (Eq. 2), a method that intentionally sharpens the aggregated soft-labels. However, the original ERA must be carefully calibrated. The foundational concept of knowledge distillation [kd] is to transfer generalization ability by using teacher’s “soft targets” (referred to as soft-labels in this paper), which contain nuanced information in the relative probabilities of incorrect answers. Excessively reducing entropy turns these soft targets into near one-hot vectors, destroying this valuable information and increases the risk of overfitting.
This balance is particularly relevant in our realistic setting where the public and private datasets differ. For instance, when a model trained on CIFAR-10 [cifar] sees a “raccoon” from CIFAR-100, its soft-label reflects a nuanced understanding of shared features (e.g., similarity to “cat” and “dog”). Aggressively sharpening this into a confident “cat” prediction would incorrectly teach the model that a raccoon is a canonical cat, distorting its learned feature representation. Preserving some entropy is thus essential.
Moreover, in the early stages of training—when the model is not yet sufficiently trained and the soft-label tends to have high entropy, with all classes assigned relatively high probabilities—strong entropy reduction may remove the information associated with the correct class and instead amplify incorrect classes, thereby leading the model to learn from an incorrect label. Thus, for such high-entropy soft-labels, it is particularly important to avoid reducing the entropy too aggressively.
This practical challenge of finding a single, robust $T$ for mixed-entropy inputs is the core limitation of ERA. The averaged soft-labels exhibit widely varying entropy levels across samples (e.g., Fig. 4a versus Fig. 4b), and, as shown in Fig. 4c, ERA’s temperature $T$ is highly sensitive to this initial entropy. Specifically, for high-entropy soft-labels (Fig. 4a), an extremely low $T$ is necessary to reduce ambiguity and extract a clear signal. However, applying such a low $T$ to low-entropy soft-labels (Fig. 4b) leads to aggressive over-sharpening. This excessive reduction destroys the nuanced information contained in the relative probabilities, effectively converting the soft-label into a one-hot vector. Conversely, setting a conservative $T$ to preserve low-entropy labels fails to sharpen high-entropy ones and can even increase their entropy. Thus, finding a single value of $T$ that balances these conflicting requirements remains widely impractical.
To address this fundamental instability, we propose Enhanced ERA, defined as:
@@
where $\overline{z^{t}}=\frac{1}{K}\sum_{k=1}^{K}z_{k}^{t}$ is the averaged local soft-labels, and $\beta$ controls the aggregation sharpness. As $\beta$ increases beyond 1, the distribution is monotonically sharpened. This is formally guaranteed by the principle of majorization: we provide a proof that the output distribution for $\beta_{2}>\beta_{1}$ majorizes the distribution for $\beta_{1}$ in Appendix B.
Beyond this core stability, Enhanced ERA directly solves the impasse of the conventional ERA. As illustrated in Fig. 4d, the power-based formulation provides smooth, consistent control. Crucially, when $\beta=1$, it perfectly preserves the original entropy, regardless of whether the input is high-entropy (Fig. 4a) or low-entropy (Fig. 4b). This provides a stable “identity” baseline, allowing the server to preserve low-entropy labels (by setting $\beta\approx 1$) while still having a mechanism to gradually and predictably sharpen high-entropy labels (by setting $\beta>1$).
This intuitive stability, observed in Fig. 4, is underpinned by a formal mathematical difference between the two methods. The instability of ERA stems from its $1/T$ (inverse) relationship and its dependence on the absolute difference between probabilities. In contrast, the stability of Enhanced ERA stems from its $\beta$ (linear) relationship and its dependence on the ratio of probabilities. A formal comparative analysis, using log-probability ratios and sensitivity derivatives to prove this “instability vs. stability” dynamic, is provided in Appendix C.
(a) Dirichlet $\alpha=0.05$
(b) Dirichlet $\alpha=0.3$
Figure 5: Preliminary comparison of aggregation stability. Learning curves show Enhanced ERA (red lines) consistently outperforms the conventional ERA (gray lines) on CIFAR-10. The optimal parameters plotted are Enhanced ERA ($\beta=2.5$) vs. ERA ($T=0.1$) for strong non-IID (5a), and Enhanced ERA ($\beta=1.0$) vs. ERA ($T=0.2$) for moderate non-IID (5b).
To validate the practical impact of this theoretical stability, we conducted a preliminary experiment (Fig. 5) comparing the two aggregation mechanisms in isolation. We evaluated the performance of the DS-FL [dsfl] framework using its default ERA mechanism, and compared it directly against an identical framework where we replaced ERA with our proposed Enhanced ERA. Neither configuration used soft-label caching, and all other experimental settings were identical to the main setup described in Section IV-A. We performed a fine-grained search for both mechanisms: ERA ($T\in\{0.025,0.05,\dots,0.275,0.3\}$) and Enhanced ERA (tuning $\beta\in\{0.5,1.0,\dots,2.5,3.0\}$). The optimal hyperparameter for each was selected based on the mean server-side test accuracy over the final 10 rounds of 3000 communication rounds.
The results in Fig. 5 show Enhanced ERA significantly outperformed in both scenarios. It yielded an accuracy improvement of approximately 3 percentage points under strong non-IID conditions ($\alpha=0.05$) and approximately 4 percentage points under moderate non-IID conditions ($\alpha=0.3$). This consistent gap provides strong practical evidence for our theoretical analysis in Appendix C: while a theoretical optimum $T$ may exist for ERA, its input-sensitive instability (Fig. 4c) makes it impractically difficult to tune. Enhanced ERA’s robust, stable control (Fig. 4d), in contrast, delivers superior performance with practical tuning.
Beyond this core stability, Enhanced ERA also provides a critical theoretical advantage regarding robustness to the number of classes $N$. The denominator of ERA (Eq. 2) scales with $N$ (i.e., $\sum e^{\overline{z}_{j}/T}\geq N$), forcing $T$ to be tuned to aggressively small values for datasets with large $N$ (e.g., CIFAR-100). In contrast, the denominator of Enhanced ERA (Eq. 4) is bounded ($\sum_{j=1}^{N}(\overline{z}_{j})^{\beta}$), making its normalization term independent of $N$ and $\beta$ a robust hyperparameter across datasets of varying class sizes.
In summary, Enhanced ERA provides controlled entropy adjustments, enabling SCARLET to effectively handle different degrees of non-IID data. This paper primarily focuses on introducing the fundamental mechanisms of soft-label caching and Enhanced ERA; the introduction of this stable, robust, and mathematically-grounded aggregation mechanism, in particular, is a key contribution. While this provides a controlled adjustment for non-IID data, the dynamic tuning of $\beta$ itself remains an open research question (discussed in Section V).